Karlottalotta:

An observer (O) spots a bird flying at a 35° angle from a line drawn horizontal to its nest. If the distance from the observer (O) to the bird (B) is 21,000 ft., how far is the bird (B) from its nest (N)? Round to the nearest whole number. 12,045 feet 14,704 feet 16,980 feet 17,202 feet

Karlottalotta:

1 attachment
ZombieLover:

1 attachment
ZombieLover:

check screenshot what does is the total?

Karlottalotta:

0.57?

XioGonz:

That isn't one of her answer choices.

Karlottalotta:

I believe is choice B

XioGonz:

sin35° = 21000/x x = 21000/sin35 = appr 14,704 feet

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Karlottalotta I believe is choice B \(\color{#0cbb34}{\text{End of Quote}}\) Correct

Karlottalotta:

thank you